OpenStudy (anonymous):

WILL FAN AND MEDAL!!! The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 4 16 128 4096 Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2

OpenStudy (anonymous):

Please helpp!!!

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

@deadpool123 @Catseyeglint911

OpenStudy (anonymous):

I personally don't know how to solve for the answer, but http://openstudy.com/study#/updates/53bd8344e4b0bbb79bdbbb68 has the answer as well as an excellent explanation. Hope it help. :)

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

thanks @Catseyeglint911

OpenStudy (anonymous):

And to be honest, just by looking at the problem above, I agree with what the person had to say. And no problem. ^^

OpenStudy (campbell_st):

well i'd say the last answer... so to get the 4th term is 2* 3rd term * 2nd term 16 = 2* 4 * 2 now try it for the 5th term 128 = 2 * 16 * 4 seems to work

OpenStudy (anonymous):

okay thanks @campbell_st, so which answer would that be? it looks like the last one

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

Thank ya'll so much!!!